Dent Enterprises (DE), a US-based medical research firm, has recently identified an island off the northeastern coast of Canada that it believes contains a large population of special spiders. The venom from these spiders will be used to fight a specific type of cancer that is unique to Europe. The owners of the island have offered to sell it to DE for USD 8 million. DE would also be faced with additional initial costs of USD 12 million to engineer the equipment that will extract the venom from the spiders. DE is an environmentally aware firm and has agreed to return the spiders to the island.

The spider population would provide enough venom to last for three years. The following information has been made available to DE:

Year

One

Two

Three

Revenues

USD 26 million

USD 28 million

USD 29 million

Operating Expenses (excluding depreciation)

USD 12 million

USD 15 million

USD 18 million

Depreciation Expense

USD 6 million

USD 2 million

Interest Expense

USD 1 million

USD 1 million

USD 1 million

Tax Rate

21%

21%

21%

Dent Enterprises has a required return on assets of 10.2%. DE will invest USD 2 million at the end year during the life of the project to plant spider friendly bushes and flowers, as well as construct a bird sanctuary on the south side of the island. DE expects to sell the island at the end of year 3 for USD 4 million.

Compute the project’s NPV.
Answer #1

Ans:For calculating the NPV we need to asertain the cash flows from 0 to 3 years which will include

  • Initial Investment
  • Net cash Flow after Tax
  • NPV Calculation
  • After asertaing the cash flow we will calculate the Present Value year by year and its formula is
  • Present value=Cash flow / (1+Expected rate of return)^Year
  • year 0 Present value= $ - 200,00,000 / (1+0.102)^0= $ -20000000
  • Year 1 Present Value= $ 11530000 /(1+0.102)^1= $ 10462794.92
  • year 2 Present value= $ 9900000 / (1+0.102)^2= $8152147.06
  • year 3 Present value = $ 10320000 / (1+0.102)^3= $7711429.86
  • After getting the value of Present value for three years we will add the same and it will be our NPV i.e $63,26,371.86 and as the NPV is positive we can go with the project.
  • Its calculation is given below

Dent Enterprises REQUIRED RETURN 10.20% Intial Investment(0 years) 1 Purchase of Island $ 80,00,000.00 2 Additional Intial Co